# Multi-Agent Deep Reinforcement Learning for Zonal Ancillary Market Coupling

### v1

@@ I Introduction
A characteristic feature of electricity is the impossibility to store it in large proportions. Hence, it must be produced exactly when it is consumed, which raises a need for a real-time monitoring of the balance between production and consumption to guarantee the stability of the power grid. When imbalances occur, mechanisms are automatically triggered to restore the nominal frequency of the electric system: a change in power production is made by some production units, which are then said to provide ancillary services, in exchange of a monetary contribution. In Western Europe, many market zones are progressively switching from an automatic and mandatory involvement of qualified production units with an amount that is fixed by the system operator to a market mechanism based on daily auctions in which the market operator publishes a demand and then the producers submit offer curves that are aggregated until they meet the demand. From the point of view of a market actor, submitting a bid in such an auction leaves some space for strategic bidding, which requires some information about the market participants. This knowledge, together with the design of a relevant bidding strategy, should enable market actors to engage their production assets in the best way in ancillary services markets, which will then give them a fair idea of the revenue they can expect. Although these ancillary services markets exist in many countries, the bidding zones can be somehow associated: for instance the Austrian and German areas can share their bids, subject to some constraints and to the availability of the transmission network across both countries. Market participants might be willing to investigate the impact of this shared market structure and to see if it stabilizes their revenues over time. The main challenges to address are the characterization of the market equilibria and the understanding of their evolution when market zones get interconnected: from the perspective of the market operator, it is important to gain insight into the behaviors of the stakeholders and to gauge the opportunity to enforce market integration over multiple countries if this can, e.g., mitigate price peak events in one country thanks to interconnection. In this article, we use noncooperative game theory to model ancillary services markets. Market equilibria are computed relying on reformulations of the game and multi-agent deep reinforcement learning (MARL). Our flexible framework allows us to assess the consequences of the interconnection of market zones by dynamically changing the coupling factors. The approach can be generalized to any number of zones. Our main contributions are summarized as follows:
@@
  • we formulate a multi-zone ancillary market model as a bilevel problem, that we subsequently reformulate as a generalized Nash game with generalized potential structure. We obtain theoretical guarantees for the existence of market equilibria;
@@
  • we propose three algorithmic solutions for the market model: two exact approaches to compute an equilibrium, integrated optimization and Gauss-Seidel best-response, and a multi-agent deep reinforcement learning (MARL) simulation. The latter allows us to run a decentralized simulation of the energy market with strategic learning and non-strategic agents;
@@ I-A Related Work
There are two main directions for research on ancillary (and zonal) market design: working on the design of the market itself, studying different formulations and their features [1]; or working on the strategic bidding aspect [2], focusing more on the point of view of the energy producers. In [3], the author discusses possible ways of clearing the electricity market and selecting bids, proposing and analyzing different options for the market operator’s objective function, the settlement rules and the pricing of the products. Previous works are mostly theoretical, comparing the different possible market designs relying on noncooperative game theory. The authors of [4] instead, propose a market designed to account for the presence of wind power producers. On the same line is the two parts work of [5, 6], where the authors first discuss the design of an ancillary service focused on primary frequency response, and then simulate this design on different test cases. Another example exploring the impact of renewable based generators, is [7], discussing bidding strategies for ancillary markets with participation of battery energy storage systems. We conclude the discussion on the formulation and features of ancillary markets with [8], which explores the consequences and opportunities of implementing ancillary services at distribution grid level, considering frequency response, voltage regulation and other parameters. Regarding the computing of bidding strategies and solutions for ancillary and zonal markets, a first contribution can be found in [9], where the authors model the market as a two-players game, analyzing the strategies of two producers going one against the other, and solving the game to obtain different market equilibria. In [10] a control based, decentralized solution is proposed to create ancillary services helping with power imbalances deriving from renewable sources. The authors in [11] formulate a zonal market as a two-stage stochastic problem, which they then recast as a mixed-integer bilinear program in order to solve it. This solution is computed from an external point of view, considering all producers at once. In [12], the market is modeled as a Stackelberg game where the market operator is the leader and the producers are the followers; the solution is then obtained by recasting the problem as a mathematical program with equilibrium constraints. Finally, [13] focuses on a game theoretic model with nonlinear convex constraints for the optimization of each producer, which in turn is simulated using a linear dynamic model. The authors then analyze theoretically the existence of a variational equilibrium and propose a distributed algorithm to compute it. The use of learning algorithm for strategic bidding has become more common in the last years, an example is [14], where the authors use deep reinforcement learning to model the bidding strategy of intelligent electricity producers. Regarding ancillary markets, the application of learning based algorithm is not particularly widespread, nonetheless, there are some examples. In [15], the authors focus on ancillary markets with hydro-based producers. The simulation they propose uses discrete inputs for the algorithm and discretized bids for the learning producers, having market power. The results are not definitive, but the learning agents show sign of strategic behavior. A more advanced approach can be found in [16], where the implemented learning algorithm utilizes continuous inputs and outputs and is specifically designed to handle multi-agent settings. The producers here only bid a price, instead of a price-quantity couple as in our case, and the focus is on the grid optimization.
@@
$\mathcal{N}(z)$ Subset of producers in $z$
@@ II An Adversarial Game Formulation
The ancillary services market can be formulated as a multi-leader single follower bilevel optimization problem. At the upper level, producers (leaders) bid strategically to maximize their revenues while anticipating the market clearing. At the lower level, the market operator (follower) determines the fractions of the bids that are accepted in order to minimize the total cost it has to pay to the producers to meet the zonal demands, while satisfying zonal, interzone and bid-related constraints. Because of the opposite objectives of the producers and market operator, the problem can be interpreted as an adversarial game, also called a zero sum game.
We define $\mathcal{N}:=\{1,\ldots,N\}$ as the set of $N$ producers and $\mathcal{Z}:=\{1,...,Z\}$ as the set of zones. Each producer $n\in\mathcal{N}$ can bid up to $K_{n}\in\mathbb{N}^{\star}$ bids. Thus, the maximum number of bids is $B:=\sum_{n\in\mathcal{N}}K_{n}$. Further, each zone $z\in\mathcal{Z}$ has three input parameters: the total requested demand $D_{z}\geq 0$, the export limit $E_{z}\geq 0$ and the core portion $C_{z}\geq 0$. For all $n\in\mathcal{N}$, $z(n)\in\mathcal{Z}$ defines the zone of producer $n$ and $\mathcal{N}(z)\subseteq\mathcal{N}$ is the subset of producers in zone $z$. Each producer $n$’s decision variable is a finite sequence of bids $y_{n}:=\text{col}(\Delta_{n},\pi_{n})=\text{col}\Big{(}(\Delta_{n,1},\pi_{n,1}),\dots,(\Delta_{n,K_{n}},\pi_{n,K_{n}})\Big{)}$, where $\Delta_{n}$ is producer $n$’s finite sequence of offers containing for each bid, the offered capacity and price. The decision variables of the market operator are $x:=(x_{n,k}^{z})_{n,k,z}\in[0,1]^{B\times Z}$, where $x_{n,k}^{z}$ represents the fraction of bid activated in each zone. A summary of the definitions is in Tab. I.
@@ II-B Market Operator
Eq. (3) ensures that the total demand in each zone is met, while Eq. (5) and Eq. (6) couple the different zones. In particular, Eq. (6) ensures that there is a given amount of the demand of a specific zone satisfied by the producers in that zone, while Eq. (5) gives a cap to the amount of energy that can be exported. Finally, Eq. (4) captures the fact that for a given bid $(n,k)$, the sum of the fractions activated in each zone is normalized. Constraints (3)-(6) can be compactly written under the form
@@
and we denote $\mathcal{X}(y):=\{x\in[0,1]^{B\times Z}:\Phi(x,y)\leq 0\}$ the feasibility set of the market operator which depends on the profile of strategies of the producers $y:=(y_{n})_{n}$.
The market operator aims to minimize the total cost it has to pay to the producers. Its objective function takes the form:
@@
Zone AZone BZonal Demand AZonal Demand BExportCore PortionCore PortionP1P2P3P4P5
@@ II-C Market Clearing and KKT Reformulation
We formulate the ancillary market clearing as a pessimistic multi-leader single follower bilevel optimization problem:
@@
With this formulation, each producer optimizes its bids without considering the different zones. Such a distinction is made by the market operator when it determines the fraction of bids accepted in each zone. Further, each producer’s objective function is independent of the other producers’ strategies. Coupling between the objective functions occurs through the market operator’s strategy. In case of multiple solutions at the lower level, the pessimistic formulation enables the selection of the worst equilibrium for the producers, obtaining a more robust formulation of the market. If the $\min$ operator were to be removed in Eq. (9a), multiple solutions might be considered at the lower level (9b)-(9c).
@@
Slater’s constraint qualification holds if one of these conditions is satisfied:
@@
We now introduce a more compact notation. First we stress that, given the dependence of the feasibility set $\mathcal{X}$ on $y$, the decision variable $x_{n}$ can be expressed as an implicit function in $y$, leading to $x_{n}\equiv x_{n}(y)$. This in turn, allows us to write the objective function as an explicit function of all producers’ decision variables, i.e. $J_{n}(y_{n},y_{-n})$. Then we let $w_{n}:=(\pi_{n,k},\Delta_{n,k},x_{n,k}^{z},\lambda_{n}^{z},\mu_{n,k},\sigma_{n},\delta_{n})\in\mathcal{W}_{n}:=[\pi^{min}_{n},\pi_{n}^{max}]^{K_{n}}\times[\tilde{\Delta}_{n},\Delta_{n}^{max}]^{K_{n}}\times[0,1]^{K_{n}\cdot Z}\times[0,+\infty]^{Z}\times[0,+\infty]^{K_{n}}\times[0,+\infty]\times[0,+\infty]$. $\mathcal{W}_{n}$ is a polyhedron with dimensionality $T:=3K+Z(K+1)+2$. The set of $l_{n}$ local constraints are captured by $h_{n}:\omega_{n}\rightarrow\mathbb{R}^{l_{n}}$ and that of the $r$ shared constraints by $G:\Omega\rightarrow\mathbb{R}^{r}$, where $\Omega_{n}\supseteq\mathcal{W}_{n}$ and $\Omega:=\prod_{n=1}^{N}\Omega_{n}$. In particular $h_{n}$ and $G$ are continuously differentiable mappings. Finally, we introduce the local feasibility set for each player $n$: $\widetilde{\mathcal{W}}_{n}:=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0\}$.
@@ III Theoretical Framework
In a large part of the literature, the convexity of the agents’ optimization problems is one indispensable assumption under which noncooperative games are studied and analyzed. On the contrary, in this paper, $\Gamma$ is a generalized Nash game with side constraints, involving trilinear objective functions and nonconvex constraints thus giving rise to a nonconvex game. In order to proceed we will use results established in [17], where the authors study a class of problems with similar settings. To that purpose, we first apply the generic setting of [17]. Then, we present a relaxed version of our problem, which is more tractable for this analysis, providing results regarding the equivalence with the non-penalized version. Further, we prove existence of market equilibrium under mild assumptions, and highlight the generalized potential structure of $\Gamma$.
III-A Nonconvex Game with Side Constraints
The formulation of [17] considers a problem where the coupling constraints are treated as penalties with associated prices, in a modified objective function. This problem is then solved as an $(N+1)$-player game, where $N$ players optimize their variables and an external one optimizes the prices, which are shared by all players. In our formulation the coupling constraints are Eqs. (3)-(5)-(6)-(12d)-(12f)-(12g), while the remaining part is considered as local constraints.
With the compact notation previously introduced, we can write the problem as an $(N+1)$-player game with prices [^1] Notice that to directly extend the results of [17], we now work with the negative of the objective function and in minimization settings.:
$\forall n\in\mathcal{N},\quad\min_{w_{n}\in\widetilde{\mathcal{W}}_{n}}\left[-J_{n}(w_{n},w_{-n})+\sum_{j=1}^{r}\rho_{j}G_{j}(w)\right]$ (14)
where we call the externally optimized price, $\rho\in\mathbb{R}^{r}_{+}$. We can now recall the definition of Nash Equilibrium that is a relevant solution concept to analyze this class of games:
Definition 1 (Nash Equilibrium with Side Constraints, [17, Definition 1]).

### v2

# Nonconvex Game and Multi Agent Reinforcement Learning for Zonal Ancillary Markets

@@ I Introduction
A characteristic feature of electricity is the impossibility to store it in large proportions. Hence, it must be produced exactly when it is consumed, which raises a need for a real-time monitoring of the balance between production and consumption to guarantee the stability of the power grid. When imbalances occur, mechanisms are automatically triggered to restore the nominal frequency of the electric system: a change in power production is made by some production units, which are then said to provide ancillary services, in exchange of a monetary contribution. In Western Europe, many market zones are progressively switching from an automatic and mandatory involvement of qualified production units with an amount that is fixed by the system operator to a market mechanism based on daily auctions in which the market operator publishes a demand and then the producers submit offer curves that are aggregated until they meet the demand. From the point of view of a market actor, submitting a bid in such an auction leaves some space for strategic bidding, which requires some information about the market participants. This knowledge, together with the design of a relevant bidding strategy, should enable market actors to engage their production assets in the best way in ancillary services markets, which will then give them a fair idea of the revenue they can expect. Although these ancillary services markets exist in many countries, the bidding zones can be somehow associated: for instance the Austrian and German areas can share their bids, subject to some constraints and to the availability of the transmission network across both countries. Market participants might be willing to investigate the impact of this shared market structure and to see if it stabilizes their revenues over time. The main challenges to address are the characterization of the market equilibria and the understanding of their evolution when market zones get interconnected: from the perspective of the market operator, it is important to gain insight into the behaviors of the stakeholders and to gauge the opportunity to enforce market integration over multiple countries if this can, e.g., mitigate price peak events in one country thanks to interconnection. In this article, we use noncooperative game theory to model ancillary services markets. Market equilibria are computed relying on reformulations of the game and multi-agent deep reinforcement learning (MARL). Our flexible framework allows us to assess the consequences of the interconnection of market zones by dynamically changing the coupling factors. The approach can be generalized to any number of zones.
Our main contributions are summarized as follows:
@@
  • we formulate a multi-zone ancillary market model as a Stackelberg game, that we subsequently reformulate as a generalized Nash game with generalized potential structure. We obtain theoretical guarantees for the existence of market equilibria;
@@
  • we propose three algorithms to compute market outcomes: an integrated optimization approach and Gauss-Seidel best-response as exact methods, to compute an equilibrium ; a multi-agent deep reinforcement learning (MARL) approach, involving strategic learning and non-learning agents, that can be implemented in a decentralized fashion, reaching a stationary solution;
@@ I-A Related Work
There are two main directions for research on ancillary (and zonal) market design: working on the design of the market itself, studying different formulations and their features [1]; or working on the strategic bidding aspect [2], focusing more on the point of view of the energy producers.
In [3], the author discusses possible ways of clearing the electricity market and selecting bids, proposing different options for the market operator’s objective function, the settlement rules and the pricing of the products. Previous works are mostly theoretical, comparing the efficiency of different market designs relying on noncooperative game theory. In [4] a two-stage electricity market design is considered, where the first stage is the day-ahead market and the second is the real time one. The authors obtain proof of equilibrium existence for the model and suggest a market mechanism to reach it. On ancillary markets, the two parts work of [5, 6], first discusses the design of an ancillary service focused on primary frequency response, and then simulates this design on different test cases. Another example exploring the impact of renewable based generators is [7], discussing bidding strategies for ancillary markets with participation of battery energy storage systems. We conclude this part on the formulation and features of ancillary markets with [8], which explores the consequences and opportunities of implementing ancillary services at distribution grid level, considering frequency response, voltage regulation and other parameters.
Regarding studies of strategic bidding in electricity market, we start by mentioning [9], where the authors focus on the spot market, obtaining a Bayesian Nash equilibrium for the problem, computed in a centralized way and also considering the effect of wind based generation. In [4] a two-stage electricity market Regarding the computing of bidding strategies and solutions for ancillary and zonal markets, a first contribution can be found in [10], where the authors model the market as a two-players game, analyzing the strategies of two producers going one against the other, and solving the game to obtain different market equilibria. In [11] a control based, decentralized solution is proposed to create ancillary services helping with power imbalances deriving from renewable sources. The authors in [12] formulate a zonal market as a two-stage stochastic problem, which they then recast as a mixed-integer bilinear program in order to solve it. This solution is computed from an external point of view, considering all producers at once. In [13], the market is modeled as a Stackelberg game where the market operator is the leader and the producers are the followers; the solution is then obtained by recasting the problem as a mathematical program with equilibrium constraints.
The use of reinforcement learning (RL) as a technique to solve complex problems related to power systems and electricity markets has become more common in the last years, even though applying to real world scenarios remains complicated due to the few theoretical results. One interesting work in this direction is [14], where the authors consider constraints to obtain guarantees for the application of an RL algorithm to voltage control. Regarding RL for strategic bidding in the electricity markets we mention [15], where the authors use deep reinforcement learning to model the bidding strategy of intelligent electricity producers. Focusing specifically ancillary markets, the application of learning based algorithm is not yet widespread, nonetheless, there are some examples. In [16], the authors focus on ancillary markets with hydro-based producers. The simulation they propose uses discrete inputs for the algorithm and discretized bids for the learning producers, having market power. The results are not definitive, but the learning agents show sign of strategic behavior. A more advanced approach can be found in [17], where the implemented learning algorithm utilizes continuous inputs and outputs and is specifically designed to handle multi-agent settings. The producers here only bid a price, instead of a price-quantity couple as in our case, and the focus is on the grid optimization.
@@
$\mathcal{N}(z)$ Subset of producers in zone $z$
@@ II An Adversarial Game Formulation
The ancillary services market can be formulated as a multi-leader single follower bilevel problem. At the upper level, producers (leaders) bid strategically to maximize their revenues while anticipating the market clearing. At the lower level, the market operator (follower) determines the fractions of the bids that are accepted in order to minimize the total cost it has to pay to the producers to meet the zonal demands, while satisfying zonal, interzone and bid-related constraints. Because of the opposite objectives of the producers and market operator, the problem can be interpreted as an adversarial game, also called a zero sum game.
We define $\mathcal{N}:=\{1,\ldots,N\}$ as the set of $N$ producers and $\mathcal{Z}:=\{1,...,Z\}$ as the set of zones. Each producer $n\in\mathcal{N}$ can bid up to $K_{n}\in\mathbb{N}^{\star}$ bids. Thus, the maximum number of bids is $B:=\sum_{n\in\mathcal{N}}K_{n}$. Further, each zone $z\in\mathcal{Z}$ has three input parameters: the total requested demand $D_{z}\geq 0$, the export limit $E_{z}\geq 0$ and the core portion $C_{z}\geq 0$. For all $n\in\mathcal{N}$, $z(n)\in\mathcal{Z}$ defines the zone of producer $n$ and $\mathcal{N}(z)\subseteq\mathcal{N}$ is the subset of producers in zone $z$. Each producer $n$’s decision variable is a finite sequence of bids $y_{n}:=\text{col}(\Delta_{n},\pi_{n})=\text{col}\Big{(}(\Delta_{n,1},\pi_{n,1}),\dots,(\Delta_{n,K_{n}},\pi_{n,K_{n}})\Big{)}$, where $\Delta_{n},\pi_{n}$ are producer $n$’s finite sequences of offered capacities and prices, respectively. The decision variables of the market operator are $x:=(x_{n,k}^{z})_{n,k,z}\in[0,1]^{B\times Z}$, where $x_{n,k}^{z}$ represents the fraction of bid activated in each zone. A summary of the definitions is in Tab. I.
@@ II-B Market Operator
Eq. (3) ensures that the total demand in each zone is met, Eq. (6) ensures that there is a given amount of the demand of a specific zone satisfied by the producers in that zone, while Eq. (5) gives a cap to the amount of energy that can be exported. Finally, Eq. (4) captures the fact that for a given bid $(n,k)$, the sum of the fractions activated in each zone is normalized. Constraints (3)-(6) can be compactly written under the form
@@
and we denote $\mathcal{X}(y):=\{x\in[0,1]^{B\times Z}:\Phi(x,y)\leq 0\}$ the feasibility set of the market operator which depends on the profile of strategies of the producers $y:=(y_{n})_{n}$. The market operator aims to minimize the total cost it has to pay to the producers. Its objective function takes the form:
@@ II-C Market Clearing and KKT Reformulation
We formulate the ancillary market clearing problem as a pessimistic multi-leader single follower bilevel optimization problem:
@@
With this formulation, each producer optimizes its bids without considering the different zones. Such a distinction is made by the market operator when it determines the fraction of bids accepted in each zone. Further, each producer’s objective function is independent of the other producers’ strategies. Coupling between the objective functions occurs through the market operator’s strategy. In case of multiple solutions at the lower level, the pessimistic formulation enables the selection of the worst equilibrium for the producers, obtaining a more robust formulation of the market.
@@
Slater’s constraint qualification holds if one of these conditions is satisfied
@@
We now introduce a more compact notation. First we stress that, given the dependence of the feasibility set $\mathcal{X}$ on $y$, the decision variable $x_{n}$ can be expressed as an implicit function in $y$, leading to $x_{n}\equiv x_{n}(y)$. This in turn, allows us to write the objective function as an explicit function of all producers’ decision variables, i.e. $J_{n}(y_{n},y_{-n})$. Then we let $w_{n}:=(\pi_{n,k},\Delta_{n,k},x_{n,k}^{z},\lambda_{n}^{z},\mu_{n,k},\sigma_{n},\delta_{n})\in\mathcal{W}_{n}:=[\pi^{min}_{n},\pi_{n}^{max}]^{K_{n}}\times[\tilde{\Delta}_{n},\Delta_{n}^{max}]^{K_{n}}\times[0,1]^{K_{n}\cdot Z}\times[0,+\infty]^{Z}\times[0,+\infty]^{K_{n}}\times[0,+\infty]\times[0,+\infty]$. $\mathcal{W}_{n}$ is a polyhedron with dimensionality $23K_{n}+Z(K_{n}+1)+$ 2. The set of $l_{n}$ local constraints are captured by $h_{n}:\omega_{n}\rightarrow\mathbb{R}^{l_{n}}$ and that of the $r$ shared constraints by $G:\Omega\rightarrow\mathbb{R}^{r}$, where $\Omega_{n}\supseteq\mathcal{W}_{n}$ and $\Omega:=\prod_{n=1}^{N}\Omega_{n}$. In particular $h_{n}$ and $G$ are continuously differentiable mappings. Finally, we introduce the local feasibility set for each player $n$: $\widetilde{\mathcal{W}}_{n}:=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}(w_{n})\leq 0\}$.
@@ III Theoretical Framework
In a large part of the literature, the convexity of the agents’ optimization problems is an essential assumption under which noncooperative games are studied and analyzed. On the contrary, in this paper, $\Gamma$ is a generalized Nash game with side constraints, involving trilinear objective functions and nonconvex constraints thus giving rise to a nonconvex game. In order to proceed we will rely on results established in [18], where the authors study a class of problems with similar settings. To that purpose, we first introduce a penalized version of our problem, which is more tractable for this analysis, providing results regarding the equivalence with the non-penalized version; then we present the settings of [18], which we use to prove existence of market equilibrium under mild assumptions. Finally we highlight the potential structure of $\Gamma$.
III-A Penalized Generalized Nash Game
We consider a penalized version of our problem, which makes it more tractable and it will allow us to use mild assumptions to obtain an equilibrium existence guarantee. We add the constraints in Eqs. (3) and (6) to the objective function, which gives the function $J_{n}^{p}(w)$ and the new constraint maps $h_{n}^{p}(w_{n})$ and $G^{p}(w)$ (see Appendix B for more details). We can formulate the penalized generalized Nash game $\Gamma^{p}$ as follows:
$\forall n\in\mathcal{N},\quad$ $\min_{w_{n}\in\mathcal{W}_{n}}$ $-J_{n}^{p}(w)$ (14a)
s.t. $h_{n}^{p}(w_{n})\leq 0$ (14b)
$G^{p}(w_{n},w_{-n})\leq 0$ (14c)
Proposition 1.
If $K_{n}\geq\frac{Z}{3}$, there exists a set of penalties such that a solution of the penalized game (14) is a solution of $\Gamma$.
Thanks to this reformulation we obtain copositive Hessians matrices associated to the constraints $h_{n}^{p}(w_{n})$ and $G^{p}(w)$, which is the key point in obtaining an existence result. We discuss the form of these matrices in Appendix D.
III-B Nonconvex Game with Side Constraints
The formulation of [18] considers a problem where the coupling constraints are treated as penalties with associated prices, in a modified objective function. This problem is then solved as an $(N+1)$-player game, where $N$ players optimize their variables and an external one optimizes the prices, which are shared by all players.
With the compact notation previously introduced, we can write the problem in Eq. (14) as an $(N+1)$-player game with prices [^1] Notice that to directly extend the results of [18], we now work with the negative of the objective function and in minimization settings.:
$\forall n\in\mathcal{N},\quad\min_{w_{n}\in\widetilde{\mathcal{W}}^{p}_{n}}\left[-J_{n}^{p}(w_{n},w_{-n})+\sum_{j=1}^{r}\rho_{j}G_{j}^{p}(w)\right]$ (15)
where we call the externally optimized price, $\rho\in\mathbb{R}^{r}_{+}$ and introduce the penalized local feasibility set $\widetilde{\mathcal{W}}_{n}^{p}:=\{w_{n}\in\mathcal{W}_{n}\,|\,h_{n}^{p}(w_{n})\leq 0\}$. We can now recall the definition of Nash Equilibrium [18, Definition 1] that is a relevant solution concept to analyze this class of games:
Definition 1 (Nash Equilibrium with Side Constraints).